OpenStudy (anonymous):

the area of the landfill i 390,000m^2. The volumetric flow rate of leachate through a compacted clay liner is 4.3*10^-5 m^3/s. The hydraulic conductivity is 6.5*10^-10m/s. Assume the head of water 0.9m. Calculate the thickness of the liner in meters.

OpenStudy (anonymous):

draw a picture first?

OpenStudy (anonymous):

What physical law is in play here?

OpenStudy (anonymous):

what i did was 6.5*10^-10m/s*0.9/390,000=1.5*10^-15 4.3*10^-5/1.5*10^-15=2.86*10^10

OpenStudy (anonymous):

i know to use darcy law